SLAC PEP-II
BABAR
SLAC<->RAL
Babar logo
HEPIC E,S & H Databases PDG HEP preprints
Organization Detector Computing Physics Documentation
Personnel Glossary Sitemap Search Hypernews
Unwrap page!
Det. Search
Who's who?
Meetings
FAQ
Images
Archive
Systems
Performance
Intern. region
Vertex Tracker
Drift chamber
DIRC
Calorimeter
IFR
LST
Magnet
Electronics
Trigger
Operations
Run Coordination
Contact Experts
Shift Takers Info
Operations Manual
Electronic Logbook
Ops Hypernews
Shift Signup
Check this page for HTML 4.01 Transitional compliance with the
W3C Validator
(More checks...)

Minutes of the Trigger Operation Meeting (August 1, 2006)

Attendees:  Rainer Bartoldus, Al Eisner, Fang Fang, Arik Kreisel, Rahmat, Rahmat,  Su Dong,  Mark Tibbetts, Dan Walker

Operations

We had another incident of  TSFX damage (out of order/out of sync) storm on Jul/30 at 23:26 (see ROM log). Slightly odd this times was that it was cured after first reconfigure but another burst of out of frame errors came 5 min later which go the shifters somewhat nervous and reconfigure again  got system back running OK again. Su Dong took a look at the IFT time tick shift complaint from DQM last week. Comparing cosmic, low lumi and current high lumi data (physics data are L3 logged events) for the IFT hit tick vs IFT pattern and projected time tick for all IFT hits, it is clear that the apparent shift of the mean hits in current data is just due to the higher rate of out of time IFT background which is flat in time. The curious observation is that the background hit pattern is dominantly U=3  (forward endcap muon) . Since U=3 is not a triggerable object for the normal physics configuration, the background are expected to be not directly related to the triggering event.  However, there is a substantial blob of U=3 correlated with the events (much larger than other patterns) which is rather odd and requires some further investigation.    

EMT

Dan set the EMT FEX prescale back to 1 and ran the EMT_PRESCALE_TEST for real datataking since Monday 6pm. He examine the dead time history for EMT in coarse vertical scale and finer vertical scale, and the total BaBar dead time. No detectable deadtime effect can be seen at the time of switching back to prescale=1 at ~18:00. The baseline deadtime is truly small (we couldn't be sure if it is 0.0025% or 0.000025%).  However, everyone got somewhat curious about the spikes, given the rather large time bin. If the prescale=1 setting continues to run well, Dan will define this as PHYSICS on Wednesday (see Dan's HN update on this after the meeting). The configuration script tool for updating the prescale is still to go into release. Dan will make a systematic check of the status of EMT software packages before passing on the baton to Mark.  

DCT/DCZ

Fang made the validation for the latest online release and it was OK. She also tried the new L1DctCalOdf after teststand code move, but unfortunately there are unresolved symbols to be tracked down still.

L3

Rainer reported that the offline release 18.7.0 validation is now formally underway. However, it is still unclear the corresponding online release will be built or  we will end up with a patch to current online release. In the following discussion of the L3 tracking code, it was brought up that Dave Brown claimed rather high rate of L3 ghost tracks (which has always being there). The proposal is to activate the ghost track filter, with code already exist in L3, but needed validation. Arik will looking into trying the filter.  

BGfilter

Rainer erported an odd incident that the OPR PC processing came to a halt on Aug/1 due to the BGfilter configuration was expired. The use of expiration date was to mark the end of the run so that OPR will not be confused by cosmic data during down time, but this got caught when Run 5 was extended through August. The expiration date will probably be set to indefinite for the future. Even before 18.7.0 validation is done, we already know that one more release 18.7.0c is needed for both the L1 trigger analysis tools and the incorporation of HighMass BG filter, but given the limited change is may bypass full validation.  

Data Quality and Monitoring

Rahmat has been working on the Lumi ratio plot. Recent release not filling some weighted quantities. Receommendation is to just use the unweighted quantities. He showed the current version of the Lumi MuMu / B-counting MuMu ratio plot (title label is incorrect). The error is calculated assuming uncorrelated statistics, but this ratio clearly has large statistic overlaps (although noone knows exactly how much). Al recommended binomial statistics. Although some polishing may be good, everyone urged to get whatever he has together right away to squeeze into 18.7.0c (otherwise may not see the code in action for another 6 months).  

L1 Configuration

The L1 trigger analysis tools are in 18.7.0c, still queued for validation before the formal collection of histograms from everyone. We need to look into how to change ZPD configuration, starting with some info on the configuration object definition.

This page is created by Selina Li (selina@slac.stanford.edu); however, you are welcome to edit and/or correct above items, specially for those who attended the meeting. Thanks.